### v1

# Agentic SLMs: Hunting Down Test Smells

Abstract
Test smells can compromise the reliability of test suites and hinder software maintenance. Although several strategies exist for detecting test smells, few address their removal. Traditional methods often rely on static analysis or machine learning, requiring significant effort and expertise. This study evaluates Llama 3.2 3B, Gemma 2 9B, DeepSeek-R1 14B, and Phi 4 14B—small, open language models—for automating the detection and refactoring of test smells through agent-based workflows. We explore workflows with one, two, and four agents across 150 instances of 5 common test smell types extracted from real-world Java projects. Unlike prior approaches, ours is easily extensible to new smells via natural language definitions and generalizes to Python and Golang. All models detected nearly all test smell instances (pass@5 of 96% with four agents), with Phi 4 14B achieving the highest refactoring accuracy (pass@5 of 75.3%). Analyses were computationally inexpensive and ran efficiently on a consumer-grade hardware. Notably, Phi 4 14B with four agents performed within 5% of proprietary models such as o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental using a single agent. Multi-agent setups outperformed single-agent ones in three out of five test smell types, highlighting their potential to improve software quality with minimal developer effort. For the Assertion Roulette smell, however, a single agent performed better. To assess practical relevance, we submitted 10 pull requests with Phi 4 14B-generated code to open-source projects. Five were merged, one was rejected, and four remain under review, demonstrating the approach’s real-world applicability.
1 Introduction
Test smells are design flaws in test code that compromise reliability and hinder maintenance [1, 2]. Similar to code smells in production code, they can lead to flakiness, false positives, and false negatives, reducing the effectiveness of test suites. These issues are prevalent in both open-source and industrial settings [3]. Common test smells include undocumented assertions (Assertion Roulette), complex conditional logic (Conditional Test Logic), duplicated assertions (Duplicate Assert), and hard-coded values (Magic Numbers[3]. Such smells harm readability and make test evolution more error-prone.
Existing tools for detecting test smells typically rely on static analysis [4, 5] or machine learning [6, 5], but these approaches are often hard to adapt to new smells or other languages [5]. Recent advances in foundation models have opened new possibilities in software engineering [7, 8], yet their use in test smell refactoring—especially through collaborative agentic workflows [9] – remains underexplored.
In this work, we evaluate small, open foundation models – Llama 3.2 3B, Gemma 2 9B, DeepSeek-R1 14B, and Phi 4 14B – for automatically detecting and refactoring test smells through agentic workflows. Our method supports one, two, or four agents and is easily extensible: new smells and refactorings can be defined in natural language. We evaluate 150 instances of 5 frequent test smells from real-world Java projects (Section 3).
Results show that all models detected nearly all test smell instances (pass@5 of 96% with four agents), with Phi 4 14B achieving the highest refactoring accuracy (pass@5 of 75.3%). This performance is within 5% of proprietary models like o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental using a single agent. Multi-agent setups outperformed single-agent ones in three of five smell types, though for Assertion Roulette, a single-agent configuration proved more effective – suggesting that workflow design should consider the specific smell being addressed. To assess real-world impact, we submitted 10 pull requests with Phi 4 14B – generated refactorings to open-source projects. Five were merged, one was rejected, and four remain open, indicating early practical viability.
Preliminary results demonstrate promising generalization across programming languages – specifically Python and Golang – using a the same setup. Beyond automating the detection and refactoring of test smells, our approach also enables developers to interact with foundation models to better understand the rationale behind suggested changes. All data and code are publicly available online [10].
2 Test Smells
Test smells are recurring patterns that reduce test clarity and maintainability. Assertion Roulette [1] occurs when multiple assertions lack descriptive messages, making it hard to trace failures. It can be mitigated by adding messages or splitting checks into separate methods. Conditional Test Logic [11] arises when tests use branching or loops, which may result in unexecuted paths. This can be addressed by isolating each condition in a dedicated test method to ensure full coverage.
Duplicate Assert [3] appears when similar assertions are repeated. Developers should either split assertions into distinct tests if they cover different behaviors or consolidate them to reduce redundancy. Exception Handling is considered a smell when tests use manual try/catch instead of framework features like assertThrows. Replacing manual handling with standard mechanisms improves test readability and intent. Magic Number occurs when numeric literals are hardcoded without explanation, impairing readability. Refactoring involves replacing literals with named constants for better clarity.
3 Methodology
The primary goal of this study is to evaluate the effectiveness of an agentic approach using small, open models to detect and remove test smells. We analyzed test methods from 11 real-world open-source GitHub projects using JUnit 5 previously studied by Soares et al. [12], including janusgraph, quarkus, testcontainers-java, opengrok, jenkins, lettuce, Mindustry, data-transfer-project, Activiti, flowable-engine, and skywalking. We found that 89% of test cases in these projects contain at most 30 LOC. We limited our analysis to tests of this size. We focused on 5 common test smell types, selecting 30 real-world examples for each. We evaluate four small, open language models – Llama 3.2 3B [13], Gemma 2 9B [14], DeepSeek-R1 14B [15], and Phi 4 14B [16] – using their default configurations. All models were accessed via the Ollama platform and executed locally on a MacBook Pro M3 with 18GB of RAM (January 2025).
Agent communication is managed using the LangChain API [17], and we apply prompting techniques such as Role (Persona) and Chain-of-Thought [18] to enhance reasoning and contextual understanding. The four-agent configuration (Figure 1) distributes responsibilities as follows: Agent 1 detects potential smells; Agent 2 confirms the detection; Agent 3 performs the refactoring; and Agent 4 evaluates the result for correctness and behavior preservation. If there is disagreement, agents engage in an Evaluator-Optimizer loop [9], repeated up to three times.
Figure 1: Agentic workflow with one, two and four agents to detect and remove test smells.
Agent 1 utilizes the following prompt to detect test smells.
{mdframed}
[backgroundcolor=cyan!5, linecolor=black, linewidth=0.5pt] You are a coding assistant with many years of experience that detects test smells.
test_smell
Your goal is to determine if the provided test code exhibits the test smell “test_smell_name”.
code
Next I may give you further details.
explanation
If the test code contains test_smell_name, respond with EXACTLY “YES” on the first line and explain why. Ignore code comments. If it does not contain, say EXACTLY “NO” on the first line and explain why not.
The prompt used by Agent 2 is designed to evaluate the output generated by Agent 1.
{mdframed}
[backgroundcolor=cyan!5, linecolor=black, linewidth=0.5pt] You are a coding expert reviewing the detection of a test smell. Consider the following test smell:
test_smell
A previous agent analyzed the following test code.
code
It gave the following answer:
explanation
Your goal is to evaluate if the previous detection by another agent is correct and justified. Ignore code comments. If you do not agree, answer NO and explain what’s wrong with it and what to correct. If yes, just say YES.
To remove a test smell, Agent 3 utilizes the prompt outlined next. During the second or third iteration, it incorporates feedback from Agent 4 (agent_feedback). {mdframed}[backgroundcolor=cyan!5, linecolor=black, linewidth=0.5pt] You are a coding assistant specializing in test code analysis and refactoring, with many years of experience.
test_smell
Your task is as follows. First analyze the provided test code to resolve test smell occurrences “test_smell_name”. If there is no smell, output the original code unchanged. Second ensure the test preserves the same behavior, but is free of test_smell_name. Third output only the final refactored code, valid under JUnit 5. Finally check the refactored version does not introduce compilation errors.
Provide only the final refactored code, with no additional explanation or text.
Code to analyze:
code
Next I may provide you further details.
agent_feedback
To evaluate the code proposed by Agent 3, Agent 4 utilizes the following prompt.
{mdframed}
[backgroundcolor=cyan!5, linecolor=black, linewidth=0.5pt] You are a code reviewer specializing in JUnit 5 test smells.
test_smell
Analyze the following code.
code
Your task is to check three conditions. First check the code does not have the test smell test_smell_name. Second verify the code follows JUnit 5 specification. Finally confirms the code does not have compilation errors. If the code satisfy all conditions, respond with EXACTLY “YES” on the first line. If not, respond with EXACTLY “NO” on the first line, then explain in one or two sentences why. Let’s think step by step.
The two-agent setup follows the same workflow but omits Agents 2 and 4 – Agent 1 handles detection, and Agent 3 performs the refactoring, without validation or feedback loops. In the single-agent setup, Agent 1 analyzes the test code, identifies potential smells, and applies the corresponding refactoring.
4 Evaluation
Table I summarizes the detailed performance of the single-agent and multi-agent approaches for detecting and refactoring test smells.
Table I: Detection and refactoring performance (pass@1) of test smells by one, two, and four agents using Llama 3.2 3B (Lla), Gemma 2 9B (Gem), DeepSeek-R1 14B (DS) and Phi 4 14B across five test smell types.
One Agent Two Agents Four Agents
Detect and Refactoring Detect Refactoring Detect Refactoring
Test Smell Sub. Phi DS Gem Lla Phi DS Gem Lla Phi DS Gem Lla Phi DS Gem Lla Phi DS Gem Lla
Assertion Roulette 30 93.3% 56.7% 66.7% 20.0% 100.0% 86.7% 96.7% 100.0% 70.0% 33.3% 36.7% 26.7% 100.0% 86.7% 96.7% 100.0% 73.3% 33.3% 36.7% 26.7%
Cond. Test Logic 30 3.3% 6.7% 6.7% 0.0% 96.7% 93.3% 100.0% 100.0% 13.3% 0.0% 13.3% 0.0% 96.7% 93.3% 100.0% 100.0% 10% 3.3% 16.7% 3.3%
Duplicate Assert 30 40.0% 36.7% 23.3% 10.0% 100.0% 50.0% 100.0% 100.0% 70.0% 20.0% 56.7% 10.0% 93.3% 60.0% 100.0% 100.0% 60.0% 20.0% 60.0% 10.0%
Exception Handling 30 43.3% 13.3% 40% 36.7% 90.0% 50.0% 93.3% 100.0% 43.3% 23.3% 43.3% 30.0% 86.7% 50.0% 90.0% 100.0% 40.0% 23.3% 40.0% 36.7%
Magic Number 30 73.3% 83.3% 30% 33.3% 93.3% 100.0% 100.0% 100.0% 80.0% 90.0% 43.3% 40.0% 90.0% 100.0% 100.0% 100.0% 70.0% 90.0% 50.0% 33.3%
4.1 Detecting Test Smells
Gemma 2 9B and Phi 4 14B exhibit good detection performance across most test smells, correctly identifying 98% and 96% instances, respectively, when using two agents. Their performance remains consistent with one or four agents. However, both models fail to detect all instances of the Exception Handling test smell. For the remaining test smells, at least one of these models achieves a perfect detection score. DeepSeek-R1 14B demonstrates the weakest performance, detecting only 78% instances. In contrast, Llama 3.2 3B consistently detects all test smells across all agent configurations (one, two, or four).
We analyze not only the binary yes or no responses but also carefully review the outputs of Agents 1 and 2 to evaluate whether their explanations make sense in detecting test smells. Next, we present examples of outputs generated by Agent 1 when using four agents. For example, Agent 1 correctly identifies a Assertion Roulette test smell with the following explanation using Phi 4 14B: {mdframed}[backgroundcolor=light-gray, linecolor=black, linewidth=0.5pt] “The provided test method testTransform exhibits the test smell known as Assertion Roulette. This occurs because the method contains multiple assertions:
  • assertEquals("123", calendar.getId());
  • assertEquals("Calendar", calendar.getName());
Each of these assertions does not include an explanatory message or parameter that clarifies what each assertion is specifically verifying. Without messages, it becomes difficult to diagnose which particular condition failed when a test fails, as all failures will simply indicate Assertion Failed without context. … ” Llama 3.2 3B, Gemma 2 9B and DeepSeek-R1 14B also detect it.
Llama 3.2 3B achieved the best performance in detecting all test smells, consistently identifying every instance. While Gemma 2 9B and Phi 4 14B also performed well, they missed a few instances. However, the explanations provided by Gemma 2 9B and Phi 4 14B were more detailed and accurate compared to those generated by Llama 3.2 3B.
These outcomes are highly encouraging and indicate that even smaller, open models can reliably detect test smells when provided with succinct, precise definitions. Unlike specialized static-analysis tools [4] that must support every possible testing framework, SLMs offer a semantic approach. We only need to provide the test smell definition in natural language. It can recognize assertions written in Mockito or JUnit without requiring an exhaustive list of method signatures to be examined.
4.2 Refactoring Test Smells
The optimal configuration for Llama 3.2 3B and Gemma 2 9B involved using four agents, successfully refactoring 22% and 40.7% out of 150 instances in one attempt, respectively. DeepSeek-R1 14B achieved its best performance with a single agent, correctly refactoring 39.3% instances. In contrast, Phi 4 14B performed best with two agents, successfully refactoring 55.3% instances, demonstrating its effectiveness in a multi-agent setup.
Using single, two, or four agents with Llama 3.2 3B, Gemma 2 9B, or Phi 4 14B, correctly remove the Exception Handling test smell. They produce the code shown in Listing 1, where the try-catch block is removed, and the fail statement is replaced with an assertThrows statement.
Listing 1: Correct refactoring for removing Exception Handling using one, two and four agents with Gemma 2 9B.
@Test
public void testStackBlowOut() {
final SmallRyeConfig config = buildConfig(maps(singletonMap("foo.blowout", "\${foo.blowout}")));
assertThrows(IllegalArgumentException.class, () -> config.getValue("foo.blowout", String.class));
In some cases, using one agent with Gemma 2 9B to refactor results in incorrect code that alters the program’s behavior by adding an extra assertion when removing the Conditional Test Logic test smell. In contrast, using two or four agents produces a correct refactoring, effectively separating the logic into two methods.
We can also correctly apply a refactoring to remove Duplicate Assert test smells using two or four agents with Phi 4 14B. However, using a single agent with Phi 4 14B applies a refactoring strategy that differs from the one proposed for this smell. Specifically, it incorporates a unique message for each assertion, effectively eliminating both this test smell and the Assertion Roulette test smell.
DeepSeek-R1 14B provides significantly more details about its reasoning process compared to the other models. In some cases, it initially proposes a valid refactored code or a correct refactoring strategy during its thought process but later revises its approach, ultimately leading to an incorrect solution.
For the Assertion Roulette test smell, Phi 4 14B achieved the best result, correctly detecting and refactoring 93.3% of instances with a single agent. In contrast, all models struggled with Conditional Test Logic, where we evaluated only one of several possible refactoring strategies. For Duplicate Assert, the highest success rate (70%) was obtained by Phi 4 14B using two agents. Phi 4 14B also performed best on Exception Handling, correctly refactoring 43.3% of cases with a single agent. The best results for Magic Number came from DeepSeek-R1 14B with two agents, achieving 90% accuracy. Overall, two- or four-agent setups led to better outcomes in three of the five test smells. However, for Assertion Roulette and Exception Handling, a single agent was sufficient. Notably, there are 23 unique instances successfully refactored by at least one model – Llama 3.2 3B (6), Gemma 2 9B (14), or DeepSeek-R1 14B (6) – that Phi 4 14B with two agents failed to refactor. By combining strengths across models, we achieve correct refactorings for 70.6% instances.
The models occasionally produced incorrect refactorings. For Assertion Roulette, common issues included misplaced messages (e.g., using JUnit 4 syntax in JUnit 5) or invalid constructs (e.g., withMessage instead of assertWithMessage). In Magic Number, the model often extracted some constants but missed others, or failed to initialize newly introduced variables. For Conditional Test Logic and Duplicate Assert, errors frequently resulted from subtle logic changes or disagreements among agents regarding whether the smell had been fully resolved. In Exception Handling, some refactorings inadvertently removed critical code along with the try block. Additionally, disagreements between Agent 3 (refactoring) and Agent 4 (evaluation) occasionally led to valid transformations being rejected—an issue observed across multiple smell types.
To improve performance under the default configuration, we conducted an experiment to assess the impact of temperature [19] settings on Phi 4 14B’s ability to detect and refactor test smells. A temperature of 0.9 produced the best results for both tasks. Additionally, we allowed Phi 4 14B with four agents to make up to five attempts, rather than limiting it to just one. With pass@5 [20], it successfully detected 96% and refactored 75.3% of the test smell instances.
4.3 Feedback Loop
The feedback loop was triggered for detection primarily in cases where Agents 2 and 4 disagreed with Agents 1 and 3, respectively (Figure 1). For detecting test smells, Gemma 2 9B and Phi 4 14B utilized the feedback loop in 1 and 9 test smell instances, respectively. In 2 instances, Phi 4 14B was able to correctly identify a test smell due to the feedback loop. Llama 3.2 3B did not require the feedback loop for detecting any test smells. DeepSeek-R1 14B successfully detected 3 instances of the Duplicate Assert test smell following a feedback loop. We employed a three-iteration feedback loop, which was sufficient for Agents 1 and 2 to reach a consensus on detecting test smells. For Agents 3 and 4 to reach a consensus on refactoring test smells, only 7.3% of the cases required more than three iterations. By allowing additional iterations, we achieved consensus in all cases by the eighth iteration, except for a single instance of the Conditional Test Logic test smell.
4.4 Pull Requests
To assess the acceptance of the refactorings generated by Phi 4 14B, which achieved the best performance in our study, we submitted 10 pull requests, ensuring at least one for each test smell type across five different open-source projects. As of this writing, five of these pull requests have been accepted and integrated: two in janusgraph (addressing the Assertion Roulette and Exception Handling test smells), and one each in opengrok (Duplicate Assert), jenkins (Duplicate Assert), and lettuce (Magic Number Test). One pull request was rejected in data-transfer-project because contributors noted that the additional message in each assertion did not improve their code. Four pull requests remain open: one in janusgraph and three in testcontainers-java, all related to Duplicate Assert and Magic Number Test smells.
4.5 Proprietary and Larger Models
In March 2025, we compared Phi 4 14B with four agents and pass@5 against proprietary models—o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental—each using a single-agent setup. o1-mini achieved a refactoring performance of 66.7%, which was inferior to Phi 4 14B with four agents. Evaluating o3-mini-high on 25 test smells that o1-mini failed to refactor yielded a 40% success rate. To further assess difficult cases, we tested Gemini 2.5 Pro Experimental, the top-ranked model on LLM Arena at the time of writing, which successfully refactored 8 out of 37 test smells that Phi 4 14B (with four agents and five attempts) failed to fix, representing a 5.3% improvement. These results indicate that Phi 4 14B, when paired with agentic workflows, can rival proprietary state-of-the-art models while remaining cost-effective and executable on consumer-grade hardware.
4.6 Prompts
We refined the prompts iteratively. For instance, in the definition of Assertion Roulette, the phrase “has more than one assertion” proved to be more precise and effective than the commonly used “has multiple non-documented assertion” found in the literature [3]. Early experiments overlooked prompt minimization, resulting in suboptimal outcomes – consistent with Hsieh et al. [21]. Providing Agent 4 with the original and refactored code plus Agent 3’s full explanation overloaded the model and reduced its effectiveness. Similarly, overly detailed smell definitions negatively impacted performance. Although not formally tested, we observed that large prompts hindered reasoning. Streamlining definitions and instructing Agent 4 to respond concisely improved efficiency. These findings align with prior studies showing that excessive or irrelevant context degrades LLM performance [22]. In some cases, we found that sentence order influenced model reasoning, supporting observations from previous work [23].
4.7 Other Languages
We evaluate whether the same setup could detect and refactor test smells in other programming languages. Using Phi 4 14B with a single agent, we analyzed the popular python-oauth2 GitHub project. The model successfully removed over 10 instances of the Assertion Roulette, Exception Handling, and Magic Number test smells across multiple test cases. Furthermore, we evaluated Phi 4 14B on the testify GitHub project using Golang, demonstrating its effectiveness in eliminating the Assertion Roulette test smell. Using a single agent, the model successfully refactored the code by adding explanatory messages to each assertion. These preliminary results indicate that the proposed approach is promising for detecting and refactoring test smells across various programming languages.
4.8 Larger Test Cases
As previously noted, 89% of the test cases in the 11 open-source projects contain at most 30 lines of code (LOC). To further evaluate Phi 4 14B, which demonstrated the best performance in our study, we selected 30 test smells (five per smell type) exceeding 30 LOC from these 11 open-source projects. We employed the same four-agent setup to detect and refactor these test smells. Although the strategy successfully identified 89% of the test smells, it showed a decline in performance compared to previous results. The effectiveness in accurately refactoring test smells also diminished, with a pass@1 rate of 28%.
4.9 Threats to Validity
One potential threat to internal validity is data leakage [24], where test smells analyzed in this study may be part of the foundation models’ training data. To mitigate this, we applied Metamorphic Testing. We limit our analysis to test code with a maximum of 30 lines of code (LOC). While this constraint excludes longer test cases, a substantial number of test cases still fall within this criterion, ensuring a representative evaluation. Another factor is prompt design, which can lead to generic responses instead of targeted explanations. To minimize variability, we used uniform and concise prompts across all evaluations, which were carefully reviewed by three authors of this paper. Validating alignment between the model output and test smell definitions is also challenging. To ensure accuracy, three authors independently reviewed selected responses, verifying whether the refactored code adhered to the intended definitions. Construct validity is limited by the small number of examples for each test smell, though they are real cases from GitHub repositories.
5 Related Work
Aljedaani et al. [5] compiled a comprehensive catalog of 22 test smell detection tools, the majority of which are designed for Java, SmallTalk, C++, and Scala, and 4 refactoring tools. Soares et al. [12] conducted a mixed-methods analysis involving 485 Java projects, exploring the adoption of JUnit 5 features to improve test code quality. Wang et al. [25] propose a tool called PyNose to detect 18 types of test smells in Python. Virgínio et al. [26] propose a tool to detect test smells in Java. Lambiase et al. [27] introduce DARTS, an IntelliJ plugin that detects and refactors three test smells: Eager Test, General Fixture, and Lack of Cohesion of Test Methods. Pontillo et al. [6] proposed a machine learning (ML)-based approach to detect test smells, focusing on four specific types. Lucas et al. [28] investigated the capability of three Large Language Models to detect test smells across multiple programming languages.
We investigate the effectiveness of agentic workflows for detecting and refactoring test smells in Java test cases. Using LangChain and Ollama, our approach coordinates up to four agents to automate these tasks. Results show that Phi 4 14B, an open model, achieves strong refactoring performance, outperforming traditional single-prompt methods in robustness and flexibility. Notably, it runs efficiently on a MacBook Pro M3 with 18GB of RAM, showing that such workflows are viable on consumer-grade hardware and perform comparably to proprietary models like o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental. A key advantage of our approach is its extensibility: new smells can be added via natural language definitions. Preliminary experiments also confirm successful application to Python and Golang, highlighting its cross-language potential. Unlike conventional methods that only detect smells or suggest code changes, our framework supports interactive conversations and explanatory feedback, helping developers understand the rationale behind each refactoring and increasing trust in automated suggestions.
6 Conclusions
In this study, we evaluated small, open language models – Llama 3.2 3B, Gemma 2 9B, DeepSeek-R1 14B, and Phi 4 14B – for the automated detection and refactoring of test smells using agentic workflows with one, two, and four agents. The evaluation covered 150 instances of 5 common test smell types from real-world projects. In addition to detection, we assessed each model’s ability to perform automated refactorings. Phi 4 14B, Gemma 2 9B, and Llama 3.2 3B successfully detected nearly all test smell instances, achieving a pass@5 of 96% with four-agent setups. Phi 4 14B delivered the best refactoring performance, reaching a pass@5 of 75.3%, and performed within 5% of proprietary models like o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental using a single agent. To assess real-world applicability, we submitted 10 pull requests with code refactored by Phi 4 14B. Five were merged, one was rejected, and four remain open – demonstrating the practical utility of our approach in live development environments.
Multi-agent workflows outperformed single-agent setups in three out of five test smell types, showing their potential to enhance code quality with minimal developer effort. However, for specific smells such as Assertion Roulette, a single agent performed better, indicating that workflow design should be tailored to the smell type. For practitioners, agentic workflows provide an automated and practical solution to maintain test code quality. For researchers, this study highlights the promise of integrating foundation models into software engineering tasks, offering a scalable path for automating test smell detection and refactoring in diverse programming environments.

### v2

@@
-Agentic SLMs: Hunting Down Test Smells
+Agentic LMs: Hunting Down Test Smells
 Abstract
-Test smells can compromise the reliability of test suites and hinder software maintenance.
-Although several strategies exist for detecting test smells, few address their removal.
-Traditional methods often rely on static analysis or machine learning, requiring significant effort and expertise.
-This study evaluates Llama 3.2 3B, Gemma 2 9B, DeepSeek-R1 14B, and Phi 4 14B—small, open language models—for automating the detection and refactoring of test smells through agent-based workflows.
-We explore workflows with one, two, and four agents across 150 instances of 5 common test smell types extracted from real-world Java projects.
-Unlike prior approaches, ours is easily extensible to new smells via natural language definitions and generalizes to Python and Golang.
-All models detected nearly all test smell instances (pass@5 of 96% with four agents), with Phi 4 14B achieving the highest refactoring accuracy (pass@5 of 75.3%).
-Analyses were computationally inexpensive and ran efficiently on a consumer-grade hardware.
-Notably, Phi 4 14B with four agents performed within 5% of proprietary models such as o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental using a single agent.
-Multi-agent setups outperformed single-agent ones in three out of five test smell types, highlighting their potential to improve software quality with minimal developer effort.
-For the Assertion Roulette smell, however, a single agent performed better.
-To assess practical relevance, we submitted 10 pull requests with Phi 4 14B-generated code to open-source projects.
-Five were merged, one was rejected, and four remain under review, demonstrating the approach’s real-world applicability.
+Test smells reduce test suite reliability and complicate maintenance.
+While many methods detect test smells, few support automated removal, and most rely on static analysis or machine learning.
+This study evaluates models with relatively small parameter counts—Llama-3.2-3B, Gemma-2-9B, DeepSeek-R1-14B, and Phi-4-14B—for their ability to detect and refactor test smells using agent-based workflows.
+We assess workflows with one, two, and four agents over 150 instances of 5 common smells from real-world Java projects.
+Our approach generalizes to Python, Golang, and JavaScript.
+All models detected nearly all instances, with Phi-4-14B achieving the best refactoring accuracy (pass@5 of 75.3%).
+Phi-4-14B with four-agents performed within 5% of proprietary LLMs (single-agent).
+Multi-agent setups outperformed single-agent ones in three of five smell types, though for Assertion Roulette, one agent sufficed.
+We submitted pull requests with Phi-4-14B-generated code to open-source projects and six were merged.
 1 Introduction
-Test smells are design flaws in test code that compromise reliability and hinder maintenance [1, 2].
-Similar to code smells in production code, they can lead to flakiness, false positives, and false negatives, reducing the effectiveness of test suites.
+Test smells are design flaws in test code that compromise reliability and hinder maintenance [1].
+They can reduce the effectiveness of test suites.
 These issues are prevalent in both open-source and industrial settings [2].
-Common test smells include undocumented assertions (Assertion Roulette), complex conditional logic (Conditional Test Logic), duplicated assertions (Duplicate Assert), and hard-coded values (Magic Numbers[3].
-Such smells harm readability and make test evolution more error-prone.
-Existing tools for detecting test smells typically rely on static analysis [4, 5] or machine learning [6, 5], but these approaches are often hard to adapt to new smells or other languages [5].
-Recent advances in foundation models have opened new possibilities in software engineering [7, 8], yet their use in test smell refactoring—especially through collaborative agentic workflows [9] – remains underexplored.
-In this work, we evaluate small, open foundation models – Llama 3.2 3B, Gemma 2 9B, DeepSeek-R1 14B, and Phi 4 14B – for automatically detecting and refactoring test smells through agentic workflows.
+Common test smells include undocumented assertions (Assertion Roulette), complex conditional logic (Conditional Test Logic), duplicated assertions (Duplicate Assert), and hard-coded values (Magic Number[2].
+Existing tools for detecting test smells typically rely on static analysis [3] or machine learning [4, 3], but these approaches are often hard to adapt to new smells or other languages [3].
+Recent advances in foundation models have opened new possibilities in software engineering, yet their use in test smell refactoring – especially through collaborative agentic workflows – remains underexplored.
+In this work, we evaluate open models (Llama-3.2-3B, Gemma-2-9B, DeepSeek-R1-14B, and Phi-4-14B) for automatically detecting and refactoring test smells through agentic workflows.
 Our method supports one, two, or four agents and is easily extensible: new smells and refactorings can be defined in natural language.
 We evaluate 150 instances of 5 frequent test smells from real-world Java projects (Section 3).
-Results show that all models detected nearly all test smell instances (pass@5 of 96% with four agents), with Phi 4 14B achieving the highest refactoring accuracy (pass@5 of 75.3%).
-This performance is within 5% of proprietary models like o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental using a single agent.
-Multi-agent setups outperformed single-agent ones in three of five smell types, though for Assertion Roulette, a single-agent configuration proved more effective – suggesting that workflow design should consider the specific smell being addressed.
-To assess real-world impact, we submitted 10 pull requests with Phi 4 14B – generated refactorings to open-source projects.
-Five were merged, one was rejected, and four remain open, indicating early practical viability.
-Preliminary results demonstrate promising generalization across programming languages – specifically Python and Golang – using a the same setup.
-Beyond automating the detection and refactoring of test smells, our approach also enables developers to interact with foundation models to better understand the rationale behind suggested changes.
-All data and code are publicly available online [10].
+Results show that all models detected nearly all test smell instances (pass@5 of 96% with four-agents), with Phi-4-14B achieving the highest refactoring accuracy (pass@5 of 75.3%).
+This performance is within 5% of proprietary LLMs using a single-agent.
+Multi-agent setups outperformed single-agent ones in three of five smell types, though for Assertion Roulette, a single-agent configuration proved more effective.
+Preliminary results demonstrate promising generalization across programming languages – specifically Python, Golang, and JavaScript – using the same setup.
+Beyond automating the detection and refactoring of test smells, our approach also enables developers to interact with models to better understand the rationale behind suggested changes.
+All artifacts are publicly available online.
+[^1] https://zenodo.org/records/16740114
 2 Test Smells
 Test smells are recurring patterns that reduce test clarity and maintainability.
-Assertion Roulette [1] occurs when multiple assertions lack descriptive messages, making it hard to trace failures.
-It can be mitigated by adding messages or splitting checks into separate methods.
-Conditional Test Logic [11] arises when tests use branching or loops, which may result in unexecuted paths.
-This can be addressed by isolating each condition in a dedicated test method to ensure full coverage.
-Duplicate Assert [3] appears when similar assertions are repeated.
-Developers should either split assertions into distinct tests if they cover different behaviors or consolidate them to reduce redundancy.
-Exception Handling is considered a smell when tests use manual try/catch instead of framework features like assertThrows.
-Replacing manual handling with standard mechanisms improves test readability and intent.
-Magic Number occurs when numeric literals are hardcoded without explanation, impairing readability.
-Refactoring involves replacing literals with named constants for better clarity.
+Assertion Roulette [1] occurs when a test method contains more than one assertion statement without an explanation or message (parameter in the assertion method).
+To mitigate this smell, developers should add a message to each assertion.
+Conditional Test Logic [5] arises when a test method contains one or more control statements (i.e., conditional expression, and loop statements).
+This can be addressed by having the developer create a dedicated test method for each condition.
+Duplicate Assert [2] appears when a test method contains more than one assertion statement with the same parameters.
+To address this, developers should split assertions that test different scenarios or states into separate tests for clarity.
+Exception Handling occurs when a test method contains either a throw statement or at least a catch clause.
+To avoid this smell, use the testing framework’s features (e.g., assertThrows) instead of manually catching or throwing exceptions.
+Magic Number occurs when a test method contains an assertion with a numeric literal as an argument.
+Refactoring involves extracting and initializing all magic numbers into constants or local variables with descriptive names.
 3 Methodology
-The primary goal of this study is to evaluate the effectiveness of an agentic approach using small, open models to detect and remove test smells.
-We analyzed test methods from 11 real-world open-source GitHub projects using JUnit 5 previously studied by Soares et al.
+The primary goal of this study is to evaluate the effectiveness of an agentic approach using models to detect and remove test smells.
+We analyzed test methods from 11 real-world open-source GitHub projects using JUnit-5 previously studied by Soares et al.
 [6], including janusgraph, quarkus, testcontainers-java, opengrok, jenkins, lettuce, Mindustry, data-transfer-project, Activiti, flowable-engine, and skywalking.
-We found that 89% of test cases in these projects contain at most 30 LOC.
-We limited our analysis to tests of this size.
-We focused on 5 common test smell types, selecting 30 real-world examples for each.
-We evaluate four small, open language models – Llama 3.2 3B [13], Gemma 2 9B [14], DeepSeek-R1 14B [15], and Phi 4 14B [16] – using their default configurations.
-All models were accessed via the Ollama platform and executed locally on a MacBook Pro M3 with 18GB of RAM (January 2025).
-Agent communication is managed using the LangChain API [17], and we apply prompting techniques such as Role (Persona) and Chain-of-Thought [18] to enhance reasoning and contextual understanding.
-The four-agent configuration (Figure 1) distributes responsibilities as follows: Agent 1 detects potential smells; Agent 2 confirms the detection; Agent 3 performs the refactoring; and Agent 4 evaluates the result for correctness and behavior preservation.
-If there is disagreement, agents engage in an Evaluator-Optimizer loop [9], repeated up to three times.
+We limited our analysis to tests with up to 30 LOC, as 89% of test cases in the studied projects fall within this range.
+We focused on 5 common test smell types [2], selecting 30 real-world examples for each.
+We evaluate smaller models up to 14B parameters: Llama-3.2-3B, Gemma-2-9B, DeepSeek-R1-14B, and Phi-4-14B using their default configurations.
+All models were accessed via the Ollama platform and executed locally on a MacBook Pro M3 with 18GB of RAM (July 2025).
+Agent communication is managed using the LangChain API, and we apply prompting techniques such as Role (Persona) and Chain-of-Thought [7] to enhance reasoning and contextual understanding.
+The four-agent configuration (Figure 1) distributes responsibilities as follows: Agent1 detects potential smells; Agent2 confirms the detection; Agent3 performs the refactoring; and Agent4 verifies the correctness of the refactored code.
+If there is disagreement, agents engage in an Evaluator-Optimizer loop [^2] https://www.anthropic.com/research/building-effective-agents, repeated up to three times.
 Figure 1: Agentic workflow with one, two and four agents to detect and remove test smells.
 Agent1 utilizes the following prompt to detect test smells.
-{mdframed}
-[backgroundcolor=cyan!5, linecolor=black, linewidth=0.5pt] You are a coding assistant with many years of experience that detects test smells.
-test_smell
+The prompts use variables to represent relevant information required for each analysis.
+For instance, test_smell_definition_and_refactoring comprises two sentences: one that defines the smell and another that suggests an appropriate refactoring (see Section 2 for example sentences).
+You are a coding assistant with many years of experience that detects test smells.
+test_smell_definition_and_refactoring
 Your goal is to determine if the provided test code exhibits the test smell “test_smell_name”.
@@
 Next I may give you further details.
-explanation
+agent_feedback
 If the test code contains test_smell_name, respond with EXACTLY “YES” on the first line and explain why.
@@
 The prompt used by Agent2 is designed to evaluate the output generated by Agent1.
-{mdframed}
-[backgroundcolor=cyan!5, linecolor=black, linewidth=0.5pt] You are a coding expert reviewing the detection of a test smell. Consider the following test smell:
-test_smell
+You are a coding expert reviewing the detection of a test smell. Consider the following test smell.
+test_smell_definition_and_refactoring
 A previous agent analyzed the following test code.
 code
-It gave the following answer:
-explanation
+It gave the following answer.
+agent_answer
 Your goal is to evaluate if the previous detection by another agent is correct and justified.
@@
 To remove a test smell, Agent3 utilizes the prompt outlined next.
-During the second or third iteration, it incorporates feedback from Agent 4 (agent_feedback).
-{mdframed}[backgroundcolor=cyan!5, linecolor=black, linewidth=0.5pt] You are a coding assistant specializing in test code analysis and refactoring, with many years of experience.
-test_smell
+During the second or third iteration, it incorporates feedback from Agent4.
+You are a coding assistant specializing in test code analysis and refactoring, with many years of experience.
+test_smell_definition_and_refactoring
 Your task is as follows.
@@
 agent_feedback
-To evaluate the code proposed by Agent 3, Agent 4 utilizes the following prompt.
-{mdframed}
-[backgroundcolor=cyan!5, linecolor=black, linewidth=0.5pt] You are a code reviewer specializing in JUnit 5 test smells.
-test_smell
+To evaluate the refactored_code proposed by Agent3, Agent4 utilizes the following prompt.
+You are a code reviewer specializing in JUnit 5 test smells.
+test_smell_definition_and_refactoring
 Analyze the following code.
-code
+refactored_code
 Your task is to check three conditions.
@@
 Let’s think step by step.
-The two-agent setup follows the same workflow but omits Agents 2 and 4 – Agent 1 handles detection, and Agent 3 performs the refactoring, without validation or feedback loops.
+The two-agent setup follows the same workflow but omits Agents2,4.
+Agent1 handles detection, and Agent3 performs the refactoring, without validation or feedback loops.
 In the single-agent setup, Agent1 analyzes the test code, identifies potential smells, and applies the corresponding refactoring.
 4 Evaluation
-Table I summarizes the detailed performance of the single-agent and multi-agent approaches for detecting and refactoring test smells.
-Table I: Detection and refactoring performance (pass@1) of test smells by one, two, and four agents using Llama 3.2 3B (Lla), Gemma 2 9B (Gem), DeepSeek-R1 14B (DS) and Phi 4 14B across five test smell types.
+Table I summarizes the detailed performance (pass@1) of the single-agent and multi-agent approaches for detecting and refactoring test smells.
+Table I: Detection and refactoring performance (pass@1) of test smells by one, two, and four agents using Llama-3.2-3B (Lla), Gemma-2-9B (Gem), DeepSeek-R1-14B (DS) and Phi-4-14B (Phi).
 One Agent Two Agents Four Agents
@@
 Assertion Roulette 30 93.3% 56.7% 66.7% 20.0% 100.0% 86.7% 96.7% 100.0% 70.0% 33.3% 36.7% 26.7% 100.0% 86.7% 96.7% 100.0% 73.3% 33.3% 36.7% 26.7%
-Cond. Test Logic 30 3.3% 6.7% 6.7% 0.0% 96.7% 93.3% 100.0% 100.0% 13.3% 0.0% 13.3% 0.0% 96.7% 93.3% 100.0% 100.0% 10% 3.3% 16.7% 3.3%
+Cond. Test Logic 30 3.3% 6.7% 6.7% 0.0% 96.7% 93.3% 100.0% 100.0% 13.3% 0.0% 13.3% 0.0% 96.7% 93.3% 100.0% 100.0% 10.0% 3.3% 16.7% 3.3%
 Duplicate Assert 30 40.0% 36.7% 23.3% 10.0% 100.0% 50.0% 100.0% 100.0% 70.0% 20.0% 56.7% 10.0% 93.3% 60.0% 100.0% 100.0% 60.0% 20.0% 60.0% 10.0%
-Exception Handling 30 43.3% 13.3% 40% 36.7% 90.0% 50.0% 93.3% 100.0% 43.3% 23.3% 43.3% 30.0% 86.7% 50.0% 90.0% 100.0% 40.0% 23.3% 40.0% 36.7%
-Magic Number 30 73.3% 83.3% 30% 33.3% 93.3% 100.0% 100.0% 100.0% 80.0% 90.0% 43.3% 40.0% 90.0% 100.0% 100.0% 100.0% 70.0% 90.0% 50.0% 33.3%
+Exception Handling 30 43.3% 13.3% 40.0% 36.7% 90.0% 50.0% 93.3% 100.0% 43.3% 23.3% 43.3% 30.0% 86.7% 50.0% 90.0% 100.0% 40.0% 23.3% 40.0% 36.7%
+Magic Number 30 73.3% 83.3% 30.0% 33.3% 93.3% 100.0% 100.0% 100.0% 80.0% 90.0% 43.3% 40.0% 90.0% 100.0% 100.0% 100.0% 70.0% 90.0% 50.0% 33.3%
 4.1 Detecting Test Smells
-Gemma 2 9B and Phi 4 14B exhibit good detection performance across most test smells, correctly identifying 98% and 96% instances, respectively, when using two agents.
+Gemma-2-9B and Phi-4-14B exhibit good detection performance across most test smells, correctly identifying 98% and 96% instances, respectively, when using two agents.
 Their performance remains consistent with one or four agents.
-However, both models fail to detect all instances of the Exception Handling test smell.
-For the remaining test smells, at least one of these models achieves a perfect detection score.
-DeepSeek-R1 14B demonstrates the weakest performance, detecting only 78% instances.
-In contrast, Llama 3.2 3B consistently detects all test smells across all agent configurations (one, two, or four).
-We analyze not only the binary yes or no responses but also carefully review the outputs of Agents 1 and 2 to evaluate whether their explanations make sense in detecting test smells.
-Next, we present examples of outputs generated by Agent 1 when using four agents.
-For example, Agent 1 correctly identifies a Assertion Roulette test smell with the following explanation using Phi 4 14B: {mdframed}[backgroundcolor=light-gray, linecolor=black, linewidth=0.5pt] “The provided test method testTransform exhibits the test smell known as Assertion Roulette.
-This occurs because the method contains multiple assertions:
-  • assertEquals("123", calendar.getId());
-  • assertEquals("Calendar", calendar.getName());
-Each of these assertions does not include an explanatory message or parameter that clarifies what each assertion is specifically verifying.
-Without messages, it becomes difficult to diagnose which particular condition failed when a test fails, as all failures will simply indicate Assertion Failed without context.
-… ” Llama 3.2 3B, Gemma 2 9B and DeepSeek-R1 14B also detect it.
-Llama 3.2 3B achieved the best performance in detecting all test smells, consistently identifying every instance.
-While Gemma 2 9B and Phi 4 14B also performed well, they missed a few instances.
-However, the explanations provided by Gemma 2 9B and Phi 4 14B were more detailed and accurate compared to those generated by Llama 3.2 3B.
-These outcomes are highly encouraging and indicate that even smaller, open models can reliably detect test smells when provided with succinct, precise definitions.
-Unlike specialized static-analysis tools [4] that must support every possible testing framework, SLMs offer a semantic approach.
-We only need to provide the test smell definition in natural language.
-It can recognize assertions written in Mockito or JUnit without requiring an exhaustive list of method signatures to be examined.
+For every test smell, at least one model achieves a perfect detection score.
+DeepSeek-R1-14B demonstrates the weakest performance, detecting only 78% instances.
+In contrast, Llama-3.2-3B consistently detects all test smells across all agent configurations.
+We review Agents1,2 outputs to assess the validity of their test smell explanations.
+Next, Agent1 uses Phi-4-14B to identify the Exception Handling test smell.
+The other models also detect it.
+“The test method testStackBlowOut includes a try-catch block to handle exceptions manually, which is indicative of the Exception Handling test smell.
+The code attempts to catch an IllegalArgumentException, rather than using testing framework features like assertThrows to verify that the exception is thrown as expected.” Llama-3.2-3B achieved the highest detection rate, identifying all test smell instances.
+Gemma-2-9B and Phi-4-14B missed a few but provided more detailed and accurate explanations.
+These results demonstrate that smaller models can reliably detect test smells when given concise natural language definitions.
+Unlike static-analysis tools, models take a semantic approach, recognizing frameworks like Mockito or JUnit without needing exhaustive method lists.
 4.2 Refactoring Test Smells
-The optimal configuration for Llama 3.2 3B and Gemma 2 9B involved using four agents, successfully refactoring 22% and 40.7% out of 150 instances in one attempt, respectively.
-DeepSeek-R1 14B achieved its best performance with a single agent, correctly refactoring 39.3% instances.
-In contrast, Phi 4 14B performed best with two agents, successfully refactoring 55.3% instances, demonstrating its effectiveness in a multi-agent setup.
-Using single, two, or four agents with Llama 3.2 3B, Gemma 2 9B, or Phi 4 14B, correctly remove the Exception Handling test smell.
-They produce the code shown in Listing 1, where the try-catch block is removed, and the fail statement is replaced with an assertThrows statement.
-Listing 1: Correct refactoring for removing Exception Handling using one, two and four agents with Gemma 2 9B.
+The optimal configuration for Llama-3.2-3B and Gemma-2-9B involved using four-agents, successfully refactoring 22% and 40.7% out of 150 instances in one attempt, respectively.
+DeepSeek-R1-14B achieved its best performance with a single-agent, correctly refactoring 39.3% instances.
+In contrast, Phi-4-14B performed best with two-agents, successfully refactoring 55.3% instances, demonstrating its effectiveness in a multi-agent setup.
+Processing 150 smells using four-agents took ~1h (Llama-3.2-3B, Gemma-2-9B), ~5h (Phi-4-14B), and ~8h (DeepSeek-R1-14B).
+Using single, two, or four-agents with Llama-3.2-3B, Gemma-2-9B, or Phi-4-14B, correctly remove the Exception Handling test smell.
+For instance, they produce the code shown in Listing 1, where the try-catch block is removed, and the fail statement is replaced with an assertThrows statement.
+Listing 1: Correct refactoring for removing Exception Handling using single and multi-agents with Gemma-2-9B.
 @Test
@@
 assertThrows(IllegalArgumentException.class, () -> config.getValue("foo.blowout", String.class));
-In some cases, using one agent with Gemma 2 9B to refactor results in incorrect code that alters the program’s behavior by adding an extra assertion when removing the Conditional Test Logic test smell.
-In contrast, using two or four agents produces a correct refactoring, effectively separating the logic into two methods.
-We can also correctly apply a refactoring to remove Duplicate Assert test smells using two or four agents with Phi 4 14B.
-However, using a single agent with Phi 4 14B applies a refactoring strategy that differs from the one proposed for this smell.
+In some cases, using single-agent with Gemma-2-9B to refactor results in incorrect code that alters the program’s behavior by adding an extra assertion when removing the Conditional Test Logic test smell.
+In contrast, using multi-agents produces a correct refactoring, effectively separating the logic into two methods.
+We can also correctly apply a refactoring to remove Duplicate Assert test smells using multi-agents with Phi-4-14B.
+However, using a single-agent with Phi-4-14B applies a refactoring strategy that differs from the one proposed for this smell.
 Specifically, it incorporates a unique message for each assertion, effectively eliminating both this test smell and the Assertion Roulette test smell.
-DeepSeek-R1 14B provides significantly more details about its reasoning process compared to the other models.
+DeepSeek-R1-14B provides significantly more details about its reasoning process compared to the other models.
 In some cases, it initially proposes a valid refactored code or a correct refactoring strategy during its thought process but later revises its approach, ultimately leading to an incorrect solution.
-For the Assertion Roulette test smell, Phi 4 14B achieved the best result, correctly detecting and refactoring 93.3% of instances with a single agent.
+For the Assertion Roulette test smell, Phi-4-14B achieved the best result, correctly detecting and refactoring 93.3% of instances with a single-agent.
 In contrast, all models struggled with Conditional Test Logic, where we evaluated only one of several possible refactoring strategies.
-For Duplicate Assert, the highest success rate (70%) was obtained by Phi 4 14B using two agents.
-Phi 4 14B also performed best on Exception Handling, correctly refactoring 43.3% of cases with a single agent.
-The best results for Magic Number came from DeepSeek-R1 14B with two agents, achieving 90% accuracy.
-Overall, two- or four-agent setups led to better outcomes in three of the five test smells.
-However, for Assertion Roulette and Exception Handling, a single agent was sufficient.
-Notably, there are 23 unique instances successfully refactored by at least one model – Llama 3.2 3B (6), Gemma 2 9B (14), or DeepSeek-R1 14B (6) – that Phi 4 14B with two agents failed to refactor.
-By combining strengths across models, we achieve correct refactorings for 70.6% instances.
-The models occasionally produced incorrect refactorings.
-For Assertion Roulette, common issues included misplaced messages (e.g., using JUnit 4 syntax in JUnit 5) or invalid constructs (e.g., withMessage instead of assertWithMessage).
-In Magic Number, the model often extracted some constants but missed others, or failed to initialize newly introduced variables.
-For Conditional Test Logic and Duplicate Assert, errors frequently resulted from subtle logic changes or disagreements among agents regarding whether the smell had been fully resolved.
-In Exception Handling, some refactorings inadvertently removed critical code along with the try block.
-Additionally, disagreements between Agent 3 (refactoring) and Agent 4 (evaluation) occasionally led to valid transformations being rejected—an issue observed across multiple smell types.
-To improve performance under the default configuration, we conducted an experiment to assess the impact of temperature [19] settings on Phi 4 14B’s ability to detect and refactor test smells.
-A temperature of 0.9 produced the best results for both tasks.
-Additionally, we allowed Phi 4 14B with four agents to make up to five attempts, rather than limiting it to just one.
-With pass@5 [20], it successfully detected 96% and refactored 75.3% of the test smell instances.
+For Duplicate Assert, the highest success rate (70%) was obtained by Phi-4-14B using two-agents.
+Phi-4-14B also performed best on Exception Handling, correctly refactoring 43.3% of cases with a single-agent.
+The best results for Magic Number came from DeepSeek-R1-14B with two agents, achieving 90% accuracy.
+Overall, two or four-agent setups led to better outcomes in three of the five test smells.
+However, for Assertion Roulette and Exception Handling, a single-agent was sufficient.
+Notably, there are 23 unique instances successfully refactored by at least one model – Llama-3.2-3B (6), Gemma-2-9B (14), or DeepSeek-R1-14B (6) – that Phi-4-14B with two-agents failed to refactor.
+By combining strengths across models, we achieve correct refactorings for 70.6% instances (pass@1).
+To improve performance under the default configuration, we conducted an experiment to assess the impact of temperature [8] settings on Phi-4-14B’s ability to detect and refactor test smells.
+A temperature of 0.9 yielded the best results for both tasks, slightly improving performance by 3.3%.
+4.3 Failure Types
+The ablation study revealed distinct failure patterns across agent configurations.
+In the single-agent setup, issues frequently included incomplete fixes, superficial code changes that failed to resolve the smell, and the introduction of invalid or non-existent constructs—often stemming from incorrect syntax or reliance on frameworks other than JUnit-5.
+In Magic Number, the model sometimes extracted only a subset of constants or failed to initialize new variables.
+Refactorings for Conditional Test Logic and Duplicate Assert occasionally introduced subtle logic changes.
+In Exception Handling, some transformations mistakenly removed critical code alongside the try block.
+Introducing additional agents reduced some errors but brought new challenges, including increased code divergence, missed detections that the single agent previously succeeded on, and frequent disagreements—particularly with Agent4 rejecting correct refactorings proposed by Agent3.
+Moreover, if the multi-agent setup had used JUnit-5 instead of JUnit-4 and correctly positioned the message parameter, its performance in Assertion Roulette would have matched that of the single-agent configuration.
 4.4 Feedback Loop
-The feedback loop was triggered for detection primarily in cases where Agents 2 and 4 disagreed with Agents 1 and 3, respectively (Figure 1).
-For detecting test smells, Gemma 2 9B and Phi 4 14B utilized the feedback loop in 1 and 9 test smell instances, respectively.
-In 2 instances, Phi 4 14B was able to correctly identify a test smell due to the feedback loop.
-Llama 3.2 3B did not require the feedback loop for detecting any test smells.
-DeepSeek-R1 14B successfully detected 3 instances of the Duplicate Assert test smell following a feedback loop.
-We employed a three-iteration feedback loop, which was sufficient for Agents 1 and 2 to reach a consensus on detecting test smells.
-For Agents 3 and 4 to reach a consensus on refactoring test smells, only 7.3% of the cases required more than three iterations.
+The feedback loop was triggered when Agents2,4 disagreed with Agents1,3, respectively (see Figure 1).
+For detecting test smells, Gemma-2-9B and Phi-4-14B utilized the feedback loop in 1 and 9 test smell instances, respectively.
+In 2 instances, Phi-4-14B was able to correctly identify a test smell due to the feedback loop.
+Llama-3.2-3B did not require the feedback loop for detecting any test smells.
+DeepSeek-R1-14B successfully detected 3 instances of the Duplicate Assert test smell following a feedback loop.
+We employed a three-iteration feedback loop, which was sufficient for Agents1,2 to reach a consensus on detecting the majority of test smells.
+For Agents3,4 to reach a consensus on refactoring test smells, only 7.3% of the cases required more than three iterations.
 By allowing additional iterations, we achieved consensus in all cases by the eighth iteration, except for a single instance of the Conditional Test Logic test smell.
 4.5 Pull Requests
-To assess the acceptance of the refactorings generated by Phi 4 14B, which achieved the best performance in our study, we submitted 10 pull requests, ensuring at least one for each test smell type across five different open-source projects.
-As of this writing, five of these pull requests have been accepted and integrated: two in janusgraph (addressing the Assertion Roulette and Exception Handling test smells), and one each in opengrok (Duplicate Assert), jenkins (Duplicate Assert), and lettuce (Magic Number Test).
-One pull request was rejected in data-transfer-project because contributors noted that the additional message in each assertion did not improve their code.
-Four pull requests remain open: one in janusgraph and three in testcontainers-java, all related to Duplicate Assert and Magic Number Test smells.
-4.5 Proprietary and Larger Models
-In March 2025, we compared Phi 4 14B with four agents and pass@5 against proprietary models—o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental—each using a single-agent setup.
-o1-mini achieved a refactoring performance of 66.7%, which was inferior to Phi 4 14B with four agents.
-Evaluating o3-mini-high on 25 test smells that o1-mini failed to refactor yielded a 40% success rate.
-To further assess difficult cases, we tested Gemini 2.5 Pro Experimental, the top-ranked model on LLM Arena at the time of writing, which successfully refactored 8 out of 37 test smells that Phi 4 14B (with four agents and five attempts) failed to fix, representing a 5.3% improvement.
-These results indicate that Phi 4 14B, when paired with agentic workflows, can rival proprietary state-of-the-art models while remaining cost-effective and executable on consumer-grade hardware.
+To assess the acceptance of the refactorings generated by Phi-4-14B, which achieved the best performance in our study, we submitted 10 pull requests, ensuring at least one for each test smell type across five different open-source projects.
+As of this writing, six of these pull requests have been reviewed and merged by human maintainers, who deemed them valuable: two in janusgraph (Assertion Roulette and Exception Handling), two in lettuce (Magic Number and Duplicate Assert), and one each in opengrok and jenkins (both addressing Duplicate Assert).
+Two pull request were rejected because contributors noted that the additional message in each assertion did not improve their code.
+Two pull requests remain open.
+4.6 LLMs
+In July 2025, we also run o3, Claude-4-Sonnet, and Gemini-2.5-Pro using a single-agent setup on 150 test smells.
+Claude-4-Sonnet, o3, Gemini-2.5-Pro achieved a refactoring performance (pass@1) of 59.3%, 74.7%, and 72.0%, respectively—though at significantly higher computational cost.
+We allowed Phi-4-14B with four-agents to make up to five attempts, rather than limiting it to just one.
+With pass@5 [9], it successfully detected 96% and refactored 75.3% of the test smell instances.
+These results indicate that the open Phi-4-14B, when paired with agentic workflows, can rival proprietary state-of-the-art LLMs while remaining cost-effective and executable on consumer-grade hardware.
 4.7 Prompts
@@
 These findings align with prior studies showing that excessive or irrelevant context degrades LLM performance [11].
-In some cases, we found that sentence order influenced model reasoning, supporting observations from previous work [23].
+In some cases, we found that sentence order influenced model reasoning, supporting prior work [12].
 4.8 Other Languages
-We evaluate whether the same setup could detect and refactor test smells in other programming languages.
-Using Phi 4 14B with a single agent, we analyzed the popular python-oauth2 GitHub project.
-The model successfully removed over 10 instances of the Assertion Roulette, Exception Handling, and Magic Number test smells across multiple test cases.
-Furthermore, we evaluated Phi 4 14B on the testify GitHub project using Golang, demonstrating its effectiveness in eliminating the Assertion Roulette test smell.
-Using a single agent, the model successfully refactored the code by adding explanatory messages to each assertion.
-These preliminary results indicate that the proposed approach is promising for detecting and refactoring test smells across various programming languages.
+We assess whether the same setup can detect and refactor test smells across different programming languages.
+Using a single-agent configuration, Phi-4-14B successfully refactored (pass@1) 24/30, 27/30, and 29/30 test smells in Golang, JavaScript, and Python, respectively.
+These instances contained Assertion Roulette, Duplicate Assert, and Magic Number smells in real-world projects such as kubernetes, react-beautiful-dnd, and django.
+The model achieved performance comparable to that observed with Java test cases.
+In the JavaScript setting, it occasionally generated Java code; however, when explicitly instructed to produce JavaScript test code, it responded correctly.
+JUnit-5 compliance can also be excluded from the prompts.
+These preliminary results suggest that the proposed approach is promising for detecting and refactoring test smells across a variety of programming languages.
+4.9 Detecting Other Types of Test Smells
+We evaluated Phi-4-14B on 30 types of test smells [15], using real-world examples—18 in Java and 12 in other languages including Ruby, Python, JavaScript, Smalltalk, and TTCN-3.
+A single-agent was used, focusing solely on detection, as the catalog does not define refactoring strategies for all smells.
+Phi-4-14B successfully detected 28 out of 30 test smells (pass@1), providing detailed justifications for each.
+The two initial failures were due to ambiguous definitions, which, once clarified, enabled the accurate detection of all smells.
+We also evaluated Phi-4-14B on 45 instances spanning four additional test-smell types in JavaScript.
+It misclassified only four cases (pass@1) of OverCommented Tests, due to an imprecise definition of when a test contains excessive comments.
+After clarifying the definition, it correctly detected all test smells.
 4.10 Larger Test Cases
-As previously noted, 89% of the test cases in the 11 open-source projects contain at most 30 lines of code (LOC).
-To further evaluate Phi 4 14B, which demonstrated the best performance in our study, we selected 30 test smells (five per smell type) exceeding 30 LOC from these 11 open-source projects.
+As previously noted, 89% of the test cases in the 11 open-source projects analyzed contain at most 30 lines of code (LOC).
+To further evaluate Phi-4-14B, which demonstrated the best performance in our study, we selected 30 test smells (five per smell type) exceeding 30 LOC from these 11 open-source projects.
 We employed the same four-agent setup to detect and refactor these test smells.
-Although the strategy successfully identified 89% of the test smells, it showed a decline in performance compared to previous results.
+Although the strategy successfully identified 86% of the test smells, it showed a decline in performance compared to previous results.
 The effectiveness in accurately refactoring test smells also diminished, with a pass@1 rate of 28%.
 4.11 Threats to Validity
-One potential threat to internal validity is data leakage [24], where test smells analyzed in this study may be part of the foundation models’ training data.
-To mitigate this, we applied Metamorphic Testing.
-We limit our analysis to test code with a maximum of 30 lines of code (LOC).
+One potential threat to internal validity is data leakage, where test smells analyzed in this study may be part of the foundation models’ training data.
+To mitigate this, we applied Metamorphic Testing.1 We limit our analysis to test code with a maximum of 30 LOC.
 While this constraint excludes longer test cases, a substantial number of test cases still fall within this criterion, ensuring a representative evaluation.
 Another factor is prompt design, which can lead to generic responses instead of targeted explanations.
-To minimize variability, we used uniform and concise prompts across all evaluations, which were carefully reviewed by three authors of this paper.
+To minimize variability, we used concise prompts across all evaluations, which were carefully reviewed by three authors of this paper.
 Validating alignment between the model output and test smell definitions is also challenging.
@@
 Soares et al.
-[12] conducted a mixed-methods analysis involving 485 Java projects, exploring the adoption of JUnit 5 features to improve test code quality.
+[6] conducted a mixed-methods analysis involving 485 Java projects, exploring the adoption of JUnit-5 features to improve test code quality.
 Wang et al.
 [13] propose a tool called PyNose to detect 18 types of test smells in Python.
-Virgínio et al.
-[26] propose a tool to detect test smells in Java.
 Lambiase et al.
@@
 Pontillo et al.
-[6] proposed a machine learning (ML)-based approach to detect test smells, focusing on four specific types.
+[4] proposed a machine learning-based approach to detect test smells, focusing on four specific types.
 Lucas et al.
-[28] investigated the capability of three Large Language Models to detect test smells across multiple programming languages.
+[15] investigated the capability of three LLMs to detect test smells across multiple programming languages.
 We investigate the effectiveness of agentic workflows for detecting and refactoring test smells in Java test cases.
-Using LangChain and Ollama, our approach coordinates up to four agents to automate these tasks.
-Results show that Phi 4 14B, an open model, achieves strong refactoring performance, outperforming traditional single-prompt methods in robustness and flexibility.
-Notably, it runs efficiently on a MacBook Pro M3 with 18GB of RAM, showing that such workflows are viable on consumer-grade hardware and perform comparably to proprietary models like o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental.
+Our approach coordinates up to four-agents to automate these tasks.
+Results show that Phi-4-14B achieves strong refactoring performance.
+It runs efficiently on consumer-grade hardware and perform comparably to LLMs.
 A key advantage of our approach is its extensibility: new smells can be added via natural language definitions.
-Preliminary experiments also confirm successful application to Python and Golang, highlighting its cross-language potential.
-Unlike conventional methods that only detect smells or suggest code changes, our framework supports interactive conversations and explanatory feedback, helping developers understand the rationale behind each refactoring and increasing trust in automated suggestions.
+Preliminary experiments also confirm successful application to Python, Golang, and JavaScript, highlighting its cross-language potential.
+Unlike conventional methods that only detect smells or suggest code changes, our approach supports interactive conversations and explanatory feedback.
 6 Conclusions
-In this study, we evaluated small, open language models – Llama 3.2 3B, Gemma 2 9B, DeepSeek-R1 14B, and Phi 4 14B – for the automated detection and refactoring of test smells using agentic workflows with one, two, and four agents.
+We evaluated Llama-3.2-3B, Gemma-2-9B, DeepSeek-R1-14B, and Phi-4-14B for the automated detection and refactoring of test smells using agentic workflows with one, two, and four agents.
 The evaluation covered 150 instances of 5 common test smell types from real-world projects.
-In addition to detection, we assessed each model’s ability to perform automated refactorings.
-Phi 4 14B, Gemma 2 9B, and Llama 3.2 3B successfully detected nearly all test smell instances, achieving a pass@5 of 96% with four-agent setups.
-Phi 4 14B delivered the best refactoring performance, reaching a pass@5 of 75.3%, and performed within 5% of proprietary models like o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental using a single agent.
-To assess real-world applicability, we submitted 10 pull requests with code refactored by Phi 4 14B.
-Five were merged, one was rejected, and four remain open – demonstrating the practical utility of our approach in live development environments.
+We assessed each model’s ability to perform automated refactorings.
+Phi-4-14B, Gemma-2-9B, and Llama-3.2-3B successfully detected nearly all test smell instances.
+Phi-4-14B delivered the best refactoring performance, reaching a pass@5 of 75.3%, and performed within 5% of proprietary models like o3, Claude-4-Sonnet, and Gemini-2.5-Pro using a single-agent.
+We submitted 10 pull requests with code refactored by Phi-4-14B.
+Six were merged, two was rejected, and two remain open – demonstrating the practical utility of our approach in live development environments.
 Multi-agent workflows outperformed single-agent setups in three out of five test smell types, showing their potential to enhance code quality with minimal developer effort.
-However, for specific smells such as Assertion Roulette, a single agent performed better, indicating that workflow design should be tailored to the smell type.
-For practitioners, agentic workflows provide an automated and practical solution to maintain test code quality.
-For researchers, this study highlights the promise of integrating foundation models into software engineering tasks, offering a scalable path for automating test smell detection and refactoring in diverse programming environments.

### v3

@@
 These issues are prevalent in both open-source and industrial settings [2].
-Common test smells include undocumented assertions (Assertion Roulette), complex conditional logic (Conditional Test Logic), duplicated assertions (Duplicate Assert), and hard-coded values (Magic Number[2].
+Common test smells include undocumented assertions (Assertion Roulette), complex conditional logic (Conditional Test Logic), duplicated assertions (Duplicate Assert), and hard-coded values (Magic Number) [2].
 Existing tools for detecting test smells typically rely on static analysis [3] or machine learning [4, 3], but these approaches are often hard to adapt to new smells or other languages [3].
@@
 All artifacts are publicly available online.
-[^1] https://zenodo.org/records/16740114
+[^1] https://zenodo.org/records/17285750
 2 Test Smells
@@
 If yes, just say YES.
-To remove a test smell, Agent3 utilizes the prompt outlined next.
-During the second or third iteration, it incorporates feedback from Agent4.
+To remove a test smell, Agent3 utilizes the prompt outlined next. During the second or third iteration, it incorporates feedback from Agent4.
 You are a coding assistant specializing in test code analysis and refactoring, with many years of experience.
@@
 Table I summarizes the detailed performance (pass@1) of the single-agent and multi-agent approaches for detecting and refactoring test smells.
+In our setting, pass@1 measures whether the model’s first attempt correctly detects or refactors a test smell, with correctness independently verified by two authors and cross-checked by a third.
 Table I: Detection and refactoring performance (pass@1) of test smells by one, two, and four agents using Llama-3.2-3B (Lla), Gemma-2-9B (Gem), DeepSeek-R1-14B (DS) and Phi-4-14B (Phi).
@@
 In contrast, Llama-3.2-3B consistently detects all test smells across all agent configurations.
-We review Agents1,2 outputs to assess the validity of their test smell explanations.
-Next, Agent1 uses Phi-4-14B to identify the Exception Handling test smell.
-The other models also detect it.
+We review Agents1,2 outputs to assess the validity of their test smell explanations. Next, Agent1 uses Phi-4-14B to identify the Exception Handling test smell. The other models also detect it.
 “The test method testStackBlowOut includes a try-catch block to handle exceptions manually, which is indicative of the Exception Handling test smell.
-The code attempts to catch an IllegalArgumentException, rather than using testing framework features like assertThrows to verify that the exception is thrown as expected.” Llama-3.2-3B achieved the highest detection rate, identifying all test smell instances.
+The code attempts to catch an IllegalArgumentException, rather than using testing framework features like assertThrows to verify that the exception is thrown as expected.”
+Llama-3.2-3B achieved the highest detection rate, identifying all test smell instances.
 Gemma-2-9B and Phi-4-14B missed a few but provided more detailed and accurate explanations.
@@
 DeepSeek-R1-14B achieved its best performance with a single-agent, correctly refactoring 39.3% instances.
-In contrast, Phi-4-14B performed best with two-agents, successfully refactoring 55.3% instances, demonstrating its effectiveness in a multi-agent setup.
+Phi-4-14B performed best with two-agents, successfully refactoring 55.3% instances, demonstrating its effectiveness in a multi-agent setup.
 Processing 150 smells using four-agents took ~1h (Llama-3.2-3B, Gemma-2-9B), ~5h (Phi-4-14B), and ~8h (DeepSeek-R1-14B).
@@
 In contrast, using multi-agents produces a correct refactoring, effectively separating the logic into two methods.
-We can also correctly apply a refactoring to remove Duplicate Assert test smells using multi-agents with Phi-4-14B.
+We can also correctly apply a refactoring to remove Duplicate Assert using multi-agents with Phi-4-14B.
 However, using a single-agent with Phi-4-14B applies a refactoring strategy that differs from the one proposed for this smell.
-Specifically, it incorporates a unique message for each assertion, effectively eliminating both this test smell and the Assertion Roulette test smell.
+It incorporates a unique message for each assertion, effectively eliminating both this test smell and the Assertion Roulette test smell.
 DeepSeek-R1-14B provides significantly more details about its reasoning process compared to the other models.
 In some cases, it initially proposes a valid refactored code or a correct refactoring strategy during its thought process but later revises its approach, ultimately leading to an incorrect solution.
-For the Assertion Roulette test smell, Phi-4-14B achieved the best result, correctly detecting and refactoring 93.3% of instances with a single-agent.
+For Assertion Roulette, Phi-4-14B achieved the best result, correctly detecting and refactoring 93.3% of instances with a single-agent.
 In contrast, all models struggled with Conditional Test Logic, where we evaluated only one of several possible refactoring strategies.
@@
 The ablation study revealed distinct failure patterns across agent configurations.
-In the single-agent setup, issues frequently included incomplete fixes, superficial code changes that failed to resolve the smell, and the introduction of invalid or non-existent constructs—often stemming from incorrect syntax or reliance on frameworks other than JUnit-5.
-In Magic Number, the model sometimes extracted only a subset of constants or failed to initialize new variables.
-Refactorings for Conditional Test Logic and Duplicate Assert occasionally introduced subtle logic changes.
-In Exception Handling, some transformations mistakenly removed critical code alongside the try block.
-Introducing additional agents reduced some errors but brought new challenges, including increased code divergence, missed detections that the single agent previously succeeded on, and frequent disagreements—particularly with Agent4 rejecting correct refactorings proposed by Agent3.
-Moreover, if the multi-agent setup had used JUnit-5 instead of JUnit-4 and correctly positioned the message parameter, its performance in Assertion Roulette would have matched that of the single-agent configuration.
+Although adding more agents reduced certain errors, it also introduced new challenges, such as missed detections previously handled by the single-agent approach, incomplete fixes, and cases where Agent4 rejected correct refactorings proposed by Agent3.
+In the aggregated analysis, the multi-agent setup outperformed the single-agent configuration in three test smells, matched its performance in Exception Handling, and underperformed only in Assertion Roulette.
+The instance-level analysis with Phi-4-14B identified 12 cases in which the single-agent succeeded but the multi-agent failed.
+These included missed detections in Conditional Test Logic and Magic Number; misplaced or invalid refactorings in Assertion Roulette, where the model sometimes generated JUnit-4 instead of JUnit-5 messages; partial fixes in Magic Number, where only a subset of constants was extracted or new variables were not properly initialized; and issues in Exception Handling, where critical code was inadvertently removed along with the try block.
+In addition, one Duplicate Assert case introduced a behavioral change.
 4.4 Feedback Loop
@@
 For Agents3,4 to reach a consensus on refactoring test smells, only 7.3% of the cases required more than three iterations.
-By allowing additional iterations, we achieved consensus in all cases by the eighth iteration, except for a single instance of the Conditional Test Logic test smell.
+By allowing additional iterations, we achieved consensus in all cases by the eighth iteration, except for a single instance of Conditional Test Logic.
 4.5 Pull Requests
-To assess the acceptance of the refactorings generated by Phi-4-14B, which achieved the best performance in our study, we submitted 10 pull requests, ensuring at least one for each test smell type across five different open-source projects.
+To assess the acceptance of the refactorings generated by Phi-4-14B, which achieved the best performance in our study, we submitted 10 pull requests, across five different open-source projects.
 As of this writing, six of these pull requests have been reviewed and merged by human maintainers, who deemed them valuable: two in janusgraph (Assertion Roulette and Exception Handling), two in lettuce (Magic Number and Duplicate Assert), and one each in opengrok and jenkins (both addressing Duplicate Assert).
-Two pull request were rejected because contributors noted that the additional message in each assertion did not improve their code.
+Two pull requests were rejected because contributors noted that the additional message in each assertion did not improve their code.
 Two pull requests remain open.
@@
 We assess whether the same setup can detect and refactor test smells across different programming languages.
-Using a single-agent configuration, Phi-4-14B successfully refactored (pass@1) 24/30, 27/30, and 29/30 test smells in Golang, JavaScript, and Python, respectively.
+Using a single-agent configuration, Phi-4-14B successfully refactored (pass@1) 24/30, 25/30, and 29/30 test smells in Golang, JavaScript, and Python, respectively.
 These instances contained Assertion Roulette, Duplicate Assert, and Magic Number smells in real-world projects such as kubernetes, react-beautiful-dnd, and django.
@@
 In the JavaScript setting, it occasionally generated Java code; however, when explicitly instructed to produce JavaScript test code, it responded correctly.
-JUnit-5 compliance can also be excluded from the prompts.
 These preliminary results suggest that the proposed approach is promising for detecting and refactoring test smells across a variety of programming languages.
@@
 4.10 Larger Test Cases
-As previously noted, 89% of the test cases in the 11 open-source projects analyzed contain at most 30 lines of code (LOC).
-To further evaluate Phi-4-14B, which demonstrated the best performance in our study, we selected 30 test smells (five per smell type) exceeding 30 LOC from these 11 open-source projects.
+As previously noted, 89% of the test cases in the 11 open-source projects analyzed contain at most 30 LOC.
+To further evaluate Phi-4-14B, which demonstrated the best performance in our study, we selected 50 test smells (ten per smell type) exceeding 30 LOC from these 11 open-source projects.
 We employed the same four-agent setup to detect and refactor these test smells.
@@
 The effectiveness in accurately refactoring test smells also diminished, with a pass@1 rate of 28%.
+4.11 Code Smells
+As a feasibility study, we evaluated Phi-4-14B on 30 Java code smells drawn from real-world projects.
+By making minor adjustments to the prompts—for example, replacing references to “test” with “code”—Phi-4-14B successfully detected all instances and correctly refactored a subset of them.
+These results indicate that the approach is promising for the automated detection and refactoring of code smells.
+We leave a deeper investigation to future work.
 4.12 Threats to Validity
@@
 Phi-4-14B delivered the best refactoring performance, reaching a pass@5 of 75.3%, and performed within 5% of proprietary models like o3, Claude-4-Sonnet, and Gemini-2.5-Pro using a single-agent.
-We submitted 10 pull requests with code refactored by Phi-4-14B.
-Six were merged, two was rejected, and two remain open – demonstrating the practical utility of our approach in live development environments.
-Multi-agent workflows outperformed single-agent setups in three out of five test smell types, showing their potential to enhance code quality with minimal developer effort.
+We submitted 10 Phi-4-14B-refactored pull requests, and six were merged, showing practical utility.
+Multi-agent workflows outperformed single-agent setups in three out of five test smell types.
+Acknowledgments
+We thank the reviewers for their feedback. This work was partially supported by CNPq (403719/2024-0, 310313/2022-8, 404825/2023-0, 443393/2023-0, 312195/2021-4), FAPESQ-PB (268/2025), and FAPESB.